element of a 500w electric iron is designed for use on a 200v supply. what value of resistance is needed to be connected in series inorder that iron can be operated from a 240v supply?
1
Expert's answer
2012-09-18T10:59:04-0400
P=U^2/R =>R=U^2/P=80 om
P,R - const, so U must be constant too
Items in
series, so I1=I2 => U1/R1=U2/R2 =>U1/U2=R1/R2 => (U=const)
200/(240-200)=200/40=4=80/R2 => R2=20om